### v2

# Explaining Sources of Uncertainty in Automated Fact-Checking

Abstract
Human-AI collaboration in knowledge-intensive tasks such as fact-checking requires understanding model uncertainty in multi‑document reasoning amid conflicting/agreeing evidence. Yet, existing methods only express uncertainty as numbers or hedges without revealing which evidence conflicts cause the uncertainty, leaving users unable to resolve disagreements. We present CLUE (Conflict-&Agreement-aware Language-model Uncertainty Explanations), a plug‑and‑play white‑box framework that, to our knowledge, is the first to generate natural‑language explanations of model uncertainty grounded in conflicting/agreeing evidence. CLUE (i) identifies span‑level claim-evidence and inter‑evidence relations that signal conflict or agreement without supervision, and (ii) uses these relations to steer explanation generation, articulating how they drive the model’s uncertainty. Across three language models and two fact‑checking datasets, CLUE produces explanations that more faithfully track model uncertainty and better align with the model’s fact‑checking decisions than span‑agnostic explanation prompting; human raters also judge them more helpful, more informative, less redundant, and more logically consistent with the input. By explicitly tying uncertainty to evidence conflicts and agreements, CLUE supports practical fact‑checking and other tasks that require reasoning over complex, conflicting information.
Explaining Sources of Uncertainty in Automated Fact-Checking
Jingyi Sunthanks: Equal contribution.  Greta Warren11footnotemark: 1  Irina Shklovski   Isabelle Augenstein University of Copenhagen {jisu, grwa, ias, augenstein}@di.ku.dk
1 Introduction
Large Language Models (LLMs) are increasingly prevalent in supporting high-stakes tasks that involve reasoning about information reliability, such as fact-checking (Wang et al., 2024; Fontana et al., 2025). To foster effective use of such models in fact-checking tasks, these models must explain the rationale for their predictions (Atanasova et al., 2020; Kotonya and Toni, 2020).
Figure 1: An example of CLUE output for a claim with two evidence passages: model verdict and certainty, and an uncertainty explanation grounded in conflict/agreement span interactions.
However, current methods in automated fact-checking have been criticised for misaligning with the practical needs of fact-checkers, who must assess source reliability and reconcile conflicting evidence, rather than merely predicting a verdict label (Warren et al., 2025; Schlichtkrull et al., 2023). Current approaches often produce explanations to justify the verdict prediction  (Atanasova et al., 2020; Stammbach and Ash, 2020; Zeng and Gao, 2024), however, they provide no support for communicating model uncertainty or surfacing agreements and conflicts within the evidence (See Fig. 2). Methods that do express uncertainty do so via numeric scores (e.g., “I am 73% confident”) that are hard to contextualise (Zimmer, 1983; Wallsten et al., 1993; van der Waa et al., 2020; Liu et al., 2020) or via natural language expressions (e.g., “I’m not sure”) that often fail to faithfully reflect model uncertainty and can inflate perceived confidence (Steyvers et al., 2025; Yona et al., 2024; Kim et al., 2024). Therefore, existing explainable fact-checking systems usually exhibit two limitations: (1) neglecting to communicate model uncertainty with explanations and (2) failing to surface evidentiary conflicts and agreements that contribute to model uncertainty. This constitutes a fundamental methodological gap, as effective fact-checking requires precisely identifying the sources of uncertainty, for example, from conflicting evidence, to guide targeted verification Graves (2017); Micallef et al. (2022).
To bridge this gap, we propose CLUE, a pipeline that generates natural language explanations (NLEs) of model uncertainty by explicitly capturing conflicts and agreements in the input (e.g., a claim and its supporting or refuting evidence), see Fig. 1. The pipeline first identifies the salient span-level interactions that are essential to the prediction of the model through an unsupervised approach, providing an input-feature explanation that highlights key relationships between separate input segments, e.g., claim and evidence Ray Choudhury et al. (2023). Previous work has shown that these interactions are faithful to the model and plausible to humans Sun et al. (2025, 2026). CLUE then converts these signals into uncertainty-aware explanations by explicitly referring to the interactions, the conflict/agreement relations they express, and how they contribute to uncertainty regarding the verdict. CLUE does not require gold-label explanations or fine-tuning, and operates entirely at inference time.
Across three language models (§3.2) and two fact-checking datasets (§3.1), we evaluate two variants of CLUE, which leverage distinct methods to guide NLE generation based on the extracted key conflicts/agreements (§2.4). Automatic evaluation shows that both variants generate explanations that are more faithful to each model’s uncertainty and agree more closely with the gold fact-checking labels than a prompting baseline that lacks conflict-/agreement-span guidance (§4.5). Human study participants likewise judge CLUE explanations as more helpful, more informative, less redundant, and more logically consistent with the input. We also observe a trade-off between two variants of our CLUE framework, one attains higher faithfulness, the other higher plausibility, highlighting a promising avenue for future work to achieve both simultaneously (§4.5).
Figure 2: Illustrative comparison of verdict-oriented fact-checking explanations (e-FEVER (Stammbach and Ash, 2020), Explain-MT (Atanasova et al., 2020), JustiLM (Zeng and Gao, 2024)) and CLUE’s uncertainty-oriented explanations. In this comparison, CLUE is the only approach that explicitly traces predictive uncertainty to claim-evidence and inter-evidence conflict/agreement across multiple evidence passages.
2 Method
2.1 Preliminaries and Overall Framework
Our objective is to explain why a LLM is uncertain about a multi-evidence fact-checking instance by grounding that uncertainty in specific agreements or conflicts within the input.
Problem setup.
Each input instance is a triple $X=(C,E_{1},E_{2})$ consisting of a claim $C$ and two evidence pieces $E_{1},E_{2}$. In this work, we explore the scenarios in which one claim and two or three evidence pieces are presented. [^1] As our framework is based on the pairwise input part interaction, it is readily applicable to scenarios where more claims and evidence pieces are involved. For clarity, we denote their concatenation as $X=[x_{1},\dots,x_{|C|+|E_{1}|+|E_{2}|}]$. The task label comes from the set $\mathcal{Y}=\{\textsc{Supports},\textsc{Refutes},\textsc{Neutral}\}$.
Pipeline overview.
Our framework comprises three stages:
  1. 1.
    Uncertainty scoring. We compute predictive entropy from the model’s answer logits to obtain a scalar uncertainty score $u(X)$ (§ 2.2). This logit-based measure is model-agnostic.
  2. 2.
    Conflicts/Agreement extraction. We capture the agreements and conflicts most relevant to the model’s reasoning by identifying the text-span interactions between $C$, $E_{1}$, and $E_{2}$ that embody these relations (§ 2.3).
  3. 3.
    Explanation generation. The model receives the extracted spans as soft constraints and produces a natural language explanation $Y_{R}=[y_{1}^{\prime},\dots,y_{r}^{\prime}]$ along with its predicted label $\hat{y}$ to the identified interactions (§ 2.4).
Outputs.
For each instance $X$, the framework returns the predicted task label $\hat{y}\in\mathcal{Y}$; the numeric uncertainty score $u(X)$; and the textual explanation $Y_{R}=[y_{1}^{\prime},\dots,y_{r}^{\prime}]$ that grounds the source of uncertainty in the specific agreements or conflicts between $C,E_{1},E_{2}$.
2.2 Predictive Uncertainty Score Generation
To quantify model uncertainty for generating an answer label on a specific input sequence, we follow previous work (Kadavath et al., 2022; Yang et al., 2025) and calculate predictive uncertainty with entropy theory, which does not require multiple runs and is widely used in open-source models.
Specifically, we define the numeric uncertainty score $u$ as the entropy of the softmax distribution over the model’s output logits for a set of candidate answers $\mathcal{Y}=\{\textsc{Supports},\textsc{Refutes},\textsc{Neutral}\}$. For each candidate label $y_{i}\in\mathcal{Y}$:
$P(y_{i}\mid X)$ $=\frac{\exp(\mathrm{logit}(y_{i}))}{\sum_{j=1}^{|\mathcal{Y}|}\exp(\mathrm{logit}(y_{j}))}$ (1)
where $\mathrm{logit}(y_{i})$ is the model’s output logit towards candidate answer $y_{i}$ given input $X$. $P(y_{i}\mid X)$ is the confidence score of model for selecting $y_{i}$ as the final answer across all candidate answers within $\mathcal{Y}$. Finally, the model’s uncertainty towards the input sequence $X$ is:
$u(X)=-\sum_{y_{i}\in\mathcal{Y}}P(y_{i}\mid X)\log P(y_{i}\mid X)$ (2)
2.3 Conflict and Agreement Span Interaction Identification for Answer Uncertainty
To surface the conflicts and agreements that drive a model’s uncertainty, we extract and then label salient span interactions among the claim $C$ and two evidence passages, $E_{1}$ and $E_{2}$.
Span interaction extraction.
For each ordered input part pair $(F,T)\in\{(C,E_{1}),(C,E_{2}),(E_{1},E_{2})\}$, we follow previous work Ray Choudhury et al. (2023); Sun et al. (2025) to extract the important span interactions and their importance score to model’s answer by (i) identifying the most important attention head to the model’s answer prediction from its final layer, (ii) obtaining its attention matrix $\mathbf{A}\in\mathbb{R}^{(|F|+|T|)\times(|F|+|T|)}$, and (iii) symmetrizing the cross-part score $a^{\prime}_{p,q}$ for each possible token interaction $(x_{p},x_{q})$ within $(F,T)$:
$a^{\prime}_{p,q}=\tfrac{1}{2}\bigl(\mathbf{A}_{p,q}+\mathbf{A}_{q,p}\bigr),\quad x_{p}\!\in\!F,\;x_{q}\!\in\!T.$
Based on these token interactions, we treat each $a^{\prime}_{p,q}$ as an edge weight in a bipartite token graph between the two input parts $(F,T)$. Then the Louvain algorithm (Blondel et al., 2008) is applied to search for communities of tokens with dense intra-cluster and sparse inter-cluster relationships. From each detected token community, one span interaction, $(\text{span}_{w},\text{span}_{v})$, can be extracted by treating the neighboring tokens of the same input part, $F$ and $T$, respectively, as a span. By averaging the token importance score $a^{\prime}_{p,q}$ of each token interaction within $(\text{span}_{w},\text{span}_{v})$, its importance score $a_{wv}$ can be obtained by
$a_{wv}=\frac{1}{|\text{span}_{w}|\,|\text{span}_{v}|}\sum_{x_{p}\in\text{span}_{w}}\!\sum_{x_{q}\in\text{span}_{v}}\!a^{\prime}_{p,q}.$ (3)
The scored interactions for $(F,T)$ form $S^{(F,T)}=\{((\text{span}_{w},\text{span}_{v}),\,a_{wv})\}$.
Relation labeling.
To tag each span interaction as agreement, disagreement, or unrelated, we use a relation labeler $L_{\text{rel}}$ that can be instantiated with either an open-weight or closed LLM. $L_{\text{rel}}$ takes the full $(C,E_{1},E_{2})$ context together with a candidate span pair and outputs one label in $\{\textit{agree},\textit{disagree},\textit{unrelated}\}$. We implement this using an instruction prompt (see App. C.1) and report the labeling model in Sec. 3.2.
After labelling all three pairs, the complete interaction set for instance $X$ is
$S_{R}=S_{R}^{(C,E_{1})}\;\cup\;S_{R}^{(C,E_{2})}\;\cup\;S_{R}^{(E_{1},E_{2})},$ (4)
where, for example, $S_{R}^{(C,E_{1})}=\{((\text{span}_{w},\text{span}_{v}),\,a_{wv},\,r_{wv})\}$. Each element links two spans with an importance score and a relation label, thereby supplying the conflict- or agreement-span interactions used in later stages.
2.4 Uncertainty Natural Language Explanation Generation
To convert the extracted conflict- and agreement-span interactions into explanations for model uncertainty, we rely on two complementary mechanisms. (i) Instruction-driven prompting embeds the spans directly in the input so the model is instructed which segments to reference. (ii) Intrinsic attention steering guides the model’s own attention toward those same segments while it is generating the NLE. Both mechanisms use self-rationalization: the model first states its verdict $\hat{y}$ and then explains $Y_{R}$, a sequencing shown to improve faithfulness over pipeline approaches (Wiegreffe et al., 2021; Marasovic et al., 2022; Siegel et al., 2025).
Instruction-driven NLE.
For each instance $X$, we rank all labelled interactions by descending importance scores and keep the top $k=3$, denoted $S_{R}(k)$, to avoid overly long explanations. These three span pairs are slotted into a three-shot prompt (See App. G.1), which instructs the model to explain how the highlighted agreements or conflicts influence its confidence. Finally, the standard transformer decoding process outputs both the predicted label $\hat{y}$ and the accompanying explanation $Y_{R}$.
Attention steering.
Instead of explicit instructions, we can guide NLE generation by modifying attention on the fly following the attention steering method in PASTA (Zhang et al., 2024b). Starting from the same $S_{R}(k)$, we collect all token indices that fall inside any selected span,
$\mathcal{I}=\bigl\{p\;:\;(\text{span}_{w},\text{span}_{v})\!\in\!S_{R}(k),\;p\!\in\!\text{span}_{w}\cup\text{span}_{v}\bigr\}.$ (5)
For each attention head $(\ell,h)$ deemed relevant to model uncertainty, let $\mathbf{A}$ be its attention matrix. We down-weight non-target tokens by $\beta$:
$\tilde{A}_{ij}$ $=\frac{A_{ij}}{Z_{i}}\begin{cases}1&\text{if }j\in\mathcal{I},\\ \beta&\text{otherwise},\end{cases}$ (6)
$Z_{i}$ $=\sum_{j\in\mathcal{I}}A_{ij}+\beta\sum_{j\notin\mathcal{I}}A_{ij}.$ (7)
All other heads remain unchanged. Following Zhang et al. (2024b), we steer $|H|=100$ heads and set $\beta=0.01$ to balance steering efficacy and prevent degeneration; see App. B for the head-selection procedure. With the steered attention in place, the transformer generates $\hat{y}$ followed by the rationale $Y_{R}$, now naturally centered on the conflict- or agreement spans that drive its uncertainty.
3 Experimental Setup
3.1 Datasets
We select two fact-checking datasets, one specific to the health domain, HealthVer Sarrouti et al. (2021), and one closer to a real-world fact-checking scenario, DRUID Hagström et al. (2024). These datasets were chosen because they provide multiple evidence pieces per claim, making them well-suited to our goal of explaining model uncertainty arising from the inter-evidence conflicts and agreements. For experiments, we select six hundred instances that consist of a claim and multiple pieces of evidence, and a golden label $y\in\{\textsc{Supports},\textsc{Refutes},\textsc{Neutral}\}$ from each dataset. [^2] While DRUID has six fine-grained fact-checking labels, we merge the labels into the above three categories to balance the label categories.
3.2 Models
We compare three generation strategies for NLEs of model uncertainty:
  • PromptBaseline: A three-shot prompt baseline extending prior few-shot NLE work (Stammbach and Ash, 2020; Zeng and Gao, 2024; Zhao et al., 2024) by explicitly asking the model to highlight conflicting or supporting spans that shape its uncertainty (see prompt template in App. G.1).
  • CLUE-Span: The instruction-driven variant of our CLUE method where the extracted span interactions are filled into a three-shot prompt to guide the explanation generation (§2.4; App. G.2).
  • CLUE-Span+Steering: The attention steering variant of our CLUE method in which the same prompt as CLUE-Span is used. Additional attention steering is applied to instinctively guide the model’s explanation generation toward the identified spans (§2.4; App. G.2).
We run the main experiments on multiple recent open-weight LLMs of comparable scale: Qwen2.5-14B-Instruct (Qwen Team, 2024), Gemma-2-9B-IT (Gemma Team, 2024), and OLMo-2-1124-13B-Instruct (Team OLMo et al., 2025) and DeepSeek-R1-Distill-Qwen-14B (DeepSeek-AI, ) (model cards in App. D). Each backbone is run through CLUE on a single NVIDIA A100-SXM 80GB GPU; we choose these models to balance instruction-following/reasoning ability with inference efficiency (see cost report in App. E).
For span-interaction relation labeling, we instantiate $L_{\text{rel}}$ with GPT-4o (OpenAI Team, 2024) to reduce relation-label noise (see App. C.3 for error propagation analysis). This step is optional: replacing $L_{\text{rel}}$ with the open-weight Qwen2.5-14B-Instruct yields comparable downstream NLE quality with a small drop (App. C.2); we discuss this modularity and future directions in Limitations.
4 Automatic Evaluation
4.1 Faithfulness
To assess whether the NLEs produced by CLUE are faithful to the model’s uncertainty, we adapt the Correlational Counterfactual Test (CCT) Siegel et al. (2024) and propose Entropy-CCT. Entropy-CCT retains CCT’s perturb-and-correlate design, but replaces model’s answer probability-shift (e.g., TVD) with predictive entropy change to directly target uncertainty, which is the object our NLEs aim to explain (see the App. F.2 for comparison).
Following Siegel et al. (2024), for each instance $X$ we construct $n$ perturbations $X^{\prime}$ by inserting a random modifier token (adjective before a noun, or adverb before a verb; App. F.1). Let $u(X)$ denote the model’s uncertainty score defined by Eq. 2. We measure the perturbation’s impact on uncertainty using Absolute Entropy Change (AEC):
$\Delta u(X)=\left|u(X)-u(X^{\prime})\right|.$ (8)
For each perturbation $X^{\prime}$, we also record whether the inserted word appears in the generated NLE, using its presence as a proxy for importance. This yields a binary mention flag $m\in\{0,1\}$ (Siegel et al., 2024; Atanasova et al., 2023).
Let $D_{m}$ denote the set of perturbed examples where the NLE mentions the inserted word and $D_{\lnot m}$ the complementary set where it does not. We compute the point-biserial correlation $r_{\text{pb}}$ (Tate, 1954) between the continuous variable $\Delta u$ and the binary flag $m$ over all perturbations. The Entropy-CCT statistic is:
$\scriptstyle\text{CCT}_{\text{entropy}}=r_{\text{pb}}=\frac{\mathbb{E}_{m}[\Delta u]-\mathbb{E}_{\lnot m}[\Delta u]}{\mathrm{Std}(\Delta u)}\cdot\sqrt{\frac{|D_{m}|\cdot|D_{\lnot m}|}{(|D_{m}|+|D_{\lnot m}|)^{2}}}.$ (9)
Here, $\mathbb{E}_{m}[\Delta u]$ and $\mathbb{E}_{\lnot m}[\Delta u]$ are the mean AEC values for the two groups, and $\mathrm{Std}(\Delta u)$ is the standard deviation across all perturbations. Higher Entropy-CCT indicates that the NLE more often mentions tokens whose insertion induces larger changes in predictive uncertainty, and thus is more faithful to reflect the model’s uncertainty.
4.2 Span-Coverage
An uncertainty explanation should surface all information conveyed by the selected span interactions. We therefore compute Span-Coverage: the fraction of reference interactions that are explicitly mentioned in the generated NLE. Let $S_{\text{NLE}}$ be the set of span interactions extracted from the explanation, and let $S_{R}(k)$ be the reference set supplied in the prompt (See §2.4). Then
$\text{Span-Coverage}\;=\;\frac{\lvert S_{\text{NLE}}\cap S_{R}(k)\rvert}{\lvert S_{R}(k)\rvert}$. (10)
A higher value indicates the NLE covers a higher proportion of the information supplied by the extracted span interactions.
4.3 Span-Extraneous
Ideally, the explanation should mention only the provided interactions and avoid introducing extraneous information. We measure the proportion of mentioned interactions that do not belong to the reference set, denoted Span-Extraneous:
$\text{Span-Extraneous}\;=\;\frac{\lvert S_{\text{NLE}}\setminus S_{R}(k)\rvert}{\lvert S_{\text{NLE}}\rvert}$. (11)
A lower value indicates closer alignment with the intended span interactions.
4.4 Label-Explanation Entailment
We evaluate the extent to which the uncertainty explanation agrees with the model’s predicted label by formulating the task as a natural-language inference (NLI) problem. First, we convert the predicted label into a hypothesis using the template “The claim is supported by / refuted by / neutral to the evidence.” The generated explanation serves as the premise. The resulting premise–hypothesis pair is fed to a widely used off-the-shelf language-inference model, DeBERTa-v3 (He et al., 2023). The Label-Explanation Entailment (LEE) score is the proportion of examples for which the NLI model predicts entailment.
4.5 Results
Model Method HealthVer DRUID
Faith $\uparrow$ Cov. $\uparrow$ Ext. $\downarrow$ LEE. $\uparrow$ Faith. $\uparrow$ Cov. $\uparrow$ Ext. $\downarrow$ LEE. $\uparrow$
Qwen2.5-14B PromptBaseline -0.028 n/a n/a 0.74 -0.080 n/a n/a 0.60
CLUE-Span 0.006 0.33 0.68 0.75 0.089 0.20 0.38 0.78
CLUE-Span+Steering 0.033 0.44 0.53 0.80 0.102 0.28 0.20 0.77
OLMo-2-1124-13B PromptBaseline -0.100 n/a n/a 0.55 -0.130 n/a n/a 0.53
CLUE-Span 0.005 0.10 0.83 0.61 0.014 0.08 0.79 0.65
CLUE-Span+Steering 0.020 0.23 0.77 0.68 0.099 0.15 0.70 0.69
Gemma-2-9B-IT PromptBaseline -0.105 n/a n/a 0.66 -0.120 n/a n/a 0.57
CLUE-Span 0.007 0.34 0.59 0.82 0.043 0.23 0.43 0.76
CLUE-Span+Steering 0.021 0.39 0.50 0.85 0.098 0.30 0.47 0.81
Table 1: Automatic uncertainty-NLE evaluation. Higher is better for Faith.(§4.1)/Cov. (§4.2)/LEE. (§4.4) and lower is better for Ext. (§4.3). n/a indicates Cov./Ext. are undefined for PromptBaseline because no reference spans are provided. Best results per metric for each dataset-model pair are in bold.
Here, we present the results of our automatic NLE evaluation for the scenario where the input consists of one Claim and two pieces of Evidence [^3] See experiments with three pieces of evidence in App. H.1. For brevity, we refer to Qwen2.5-14B-Instruct, OLMo-2-1124-13B-Instruct, and Gemma-2-9B-it simply as Qwen, OLMo, and Gemma, respectively [^4] See experiments on the reasoning model DeepSeek‑R1‑Distill‑Qwen‑14B in App. H.2 .
Faithfulness.
We use Entropy-CCT, a point–biserial correlation $r_{\text{pb}}$ (Eq. 9) bounded by [1,1][-1,1] to measure the faithfulness of NLEs to the model’s uncertainty (§4.1). When $r_{\text{pb}}=0$, the explanation mentions high- and low-impact perturbation words equally often; every +0.01+0.01 adds roughly one percentage point (pp) to the chance that the explanation names a token that is truly influential for the model’s predictive uncertainty (App. F.3).
Table 1 shows that PromptBaseline is non-faithful in all six settings with all $r_{\text{pb}}$ values negative ranging from 0.03-0.03 to 0.13-0.13. Thus its NLEs mention truly influential tokens 3–13 pp less often than uninfluential ones-the opposite of faithful behaviour. Both variants of our CLUE reverse this trend. Presenting span interactions in the prompt (CLUE-Span) raises every correlation to non-negative values and peaks at $r_{\text{pb}}=0.089$ on the DRUID–Qwen setting. This means the explanation now mentions about 17 pp more often than PromptBaseline ($r_{\text{pb}}=-0.080$). Adding attention steering (CLUE-Span+Steering) lifts the $r_{\text{pb}}$ scores to $0.033$ on HealthVer and $0.102$ on DRUID with Qwen model, i.e., net gains of +6 pp and +18 pp over PromptBaseline. Moreover, four of the six positive correlations produced by CLUE-Span+Steering are significant at $p<0.01$ (Table 7 in App. F.3), confirming that the improvements are both substantial and statistically reliable. Particularly large jumps of OLMo on Druid dataset (up to $\Delta r_{\text{pb}}=+0.23\approx+23$ pp) suggest that span-interaction guidance from our CLUE framework is most beneficial for models that initially struggle to align explanations with its uncertainty.
Other Properties
We evaluate three further properties of the generated NLEs: (i) Span-Coverage of extracted conflict-/agreement- span interactions (§4.2), (ii) Span-Extraneous: mention of non-extracted spans (§4.3), and (iii) Label-Explanation Entailment with the generated fact-checking label (§4.4). As Table 1 shows, CLUE-Span+Steering outperforms CLUE-Span in both Span-Coverage and Span-Extraneous, consistent with the attention steering method’s effectiveness in directing the model to focus on provided spans during generation (Zhang et al., 2024b). Absolute numbers, however, remain modest (peak Span-Coverage: .44, Span-Extraneous: .20 with Qwen). A Span-Coverage of 1 means the NLE cites every extracted interaction, while a Span-Extraneous score of 0 means it adds none beyond them. This gap highlights considerable headroom for better integrating critical span interactions into the explanations. Among the three backbones, Qwen attains the highest Span-Coverage and the lowest Span-Extraneous scores, a trend that likely reflects its stronger instruction-following ability (see benchmark scores in App. A), and thus larger or more capable models might further narrow the gap. Both variants of our framework achieve stronger label-explanation entailment scores than the baseline, yielding explanations logically consistent with the predicted labels while remaining faithful to the model’s uncertainty.
5 Human Evaluation
5.1 Method
We recruited N=12 participants from Prolific (https://www.prolific.com/) to evaluate 120 explanations generated by PromptBaseline, CLUE-Span, CLUE-Span+Steering for 40 unique claims (20 from DRUID, 20 from HealthVer) (see details of participants and setup in App. I.1). Adapting Atanasova et al. (2020), participants ranked explanations in descending order (1st, 2nd, 3rd) according to five criteria, complementary to our automatic evaluation metrics:
  • Helpfulness. The explanation offers information that aids readers to fact-check the claim.
  • Coverage. The explanation captures all salient information in the input that matters for the fact check, distinct from Span-Coverage (§4.2), which counts overlap with pre-extracted spans.
  • Non-redundancy. The explanation does not offer irrelevant or repetitive information to the input, distinct from Span-Extraneous (§4.3) which counts mentions outside the extracted spans.
  • Consistency. The explanation contains logically consistent statements to the input, distinct from Label-Explanation Entailment (§4.4), which measures label-explanation alignment.
  • Overall Quality. Ranking of explanations by considering all criteria above.
5.2 Results
Our evaluation shows that explanations generated by CLUE are consistently preferred to PromptBaseline (Table 2). Similar to prior work that has evaluated NLEs Atanasova et al. (2020); Huang et al. (2024); Solano et al. (2024), annotator agreement was moderate to low (App. I.2.1), which we attribute to the complexity of the task.
PromptBaseline CLUE
Helpfulness
Overall 0.281 0.73
DRUID 0.312 0.688
HealthVer 0.25 0.772
Consistency
Overall 0.290 0.721
DRUID 0.309 0.691
HealthVer 0.27 0.751
Non-redundancy
Overall 0.252 0.762
DRUID 0.261 0.739
HealthVer 0.242 0.784
Coverage
Overall 0.275 0.722
DRUID 0.283 0.717
HealthVer 0.266 0.727
Overall Quality
Overall 0.325 0.678
DRUID 0.313 0.688
HealthVer 0.336 0.667
Table 2: Proportion of times explanations were ranked as participants’ first preference for each evaluation metric. Explanations generated by Qwen2.5-14B-Instruct (chosen for their high faithfulness; see §4.5)
The explanations generated by our CLUE framework are ranked as participants’ first preference 67-78% of the time, compared to 24-34% of the time for PromptBaseline. The explanations generated by CLUE are rated as most helpful, containing the least amount of redundant information, highest coverage, consistency, and overall quality approximately twice as often as those generated using PromptBaseline.
Although both CLUE-Span and CLUE-Span+Steering outperform PromptBaseline in automatic metrics (§4.5) and are ranked higher than PromptBaseline in the human study (Table 11 in App. I.2.2), we observe a mild faithfulness-plausibility trade-off when comparing the two CLUE variants. CLUE-Span+Steering achieves the highest automatic faithfulness, whereas CLUE-Span is ranked slightly higher in human-perceived overall quality. A possible reason is that CLUE-Span+Steering adheres closely to the top-k=3k{=}3 span interactions (higher Span-Coverage and lower Span-Extraneous), which improves faithfulness but can reduce fluency/naturalness when some extracted spans are fragmentary. In contrast, CLUE-Span may capture additional points that participants deemed important, but not captured by the extracted span interactions (Ray Choudhury et al., 2023). This pattern aligns with the well-documented trade-off between faithfulness and plausibility (Agarwal et al., 2024; Atanasova et al., 2023; Lu and Ma, 2024), and calls for future work to enhance both aspects for improved NLE generation, e.g., by improving the quality of the extracted span interactions.
6 Related Work
6.1 Uncertainty Quantification in LLMs
Recent work estimates LLM uncertainty primarily with white-box, logit-derived measures: predictive entropy of answer distributions (Kadavath et al., 2022; Yang et al., 2025), aggregation across generations (Malinin and Gales, 2021), and response consistency via semantic similarity (Duan et al., 2024; Kuhn et al., 2023; Nikitin et al., 2024). For closed-source models, uncertainty is typically elicited through verbalised confidence (Lin et al., 2022; Mielke et al., 2022), which is overconfident and unreliable (Yona et al., 2024; Tanneru et al., 2024), or approximated by diversity under paraphrased prompts (Zhang et al., 2024a; Chen and Mueller, 2024), which is compute-intensive and conflates prompt noise with model uncertainty. Accordingly, our method focuses on open-source models and adopts predictive entropy, a simple, interpretable, efficient white-box metric computed from answer logits, avoiding prompt-induced noise.
6.2 Linguistic Expressions of Uncertainty
Numerical uncertainty estimates are difficult for end-users to act upon as they do not address sources of uncertainty (Warren et al., 2025). Linguistic expressions of uncertainty, e.g., phrases such as “I’m sure” (Mielke et al., 2022; Tian et al., 2023; Xiong et al., 2024; Ji et al., 2025; Farquhar et al., 2024; Kim et al., 2024), may be more intuitive to understand than numerical ones (Zimmer, 1983; Wallsten et al., 1993; Windschitl and Wells, 1996). However, they are not necessarily faithful reflections of the model’s uncertainty (Yona et al., 2024; Tanneru et al., 2024) risking misleading users (Steyvers et al., 2025). Moreover, they do not explain why the model is uncertain, as our CLUE method does.
6.3 Generating Natural Language Explanations for Fact-Checking
Natural language explanations (NLEs) justify model predictions for lay readers in fact-checking (Wei Jie et al., 2024), complementing summarization and social-media work that provides graphs, manipulated spans as rationales (Ribeiro et al., 2022; Chan et al., 2023; Huang et al., 2025; Sun et al., 2026). In fact-checking, early systems extracted key sentences from supplied articles as explanations (Atanasova et al., 2020); later work improved NLE fluency via post-editing (Jolly et al., 2022) and on data from fact-checking websites (Feher et al., 2025; Raffel et al., 2020; Beltagy et al., 2020). More recent methods use highlight-based explanations as signals to enhance NLE quality, employing techniques such as graph-based modeling (Yuan et al., 2025) and self-refinement (Wang and Atanasova, 2025). Another line of work focuses on few-shot approaches that avoid model supervision, including prompting GPT-3 to produce evidence summaries (Brown et al., 2020; Stammbach and Ash, 2020), inserting a planning step (Zhao et al., 2024), and leveraging retrieval-augmented language models (Zeng and Gao, 2024). However, existing methods are often either unfaithful to model processes (Atanasova et al., 2023; Siegel et al., 2024, 2025) or fail to address model uncertainty (Warren et al., 2025), limiting their practical utility in fact-checking (Schmitt et al., 2024). Recent work in a fact-checking context has shown that explanations of model uncertainty were judged to be more helpful than providing the model verdict and numerical uncertainty alone (Warren et al., 2026), suggesting that generating such explanations may assist people to reason about the reliability of automated fact-checking predictions. Our framework addresses these gaps by explicitly explaining sources of uncertainty in automatic fact-checking, yielding explanations that are both faithful to model uncertainty and provide actionable insights for fact-checkers.
7 Conclusion
We present the first framework, CLUE, for generating NLEs of model uncertainty by referring to the conflicts and agreements between claims and multiple pieces of evidence in a fact-checking task. Our method, evaluated across three language models and two datasets, demonstrates significant improvements in both faithfulness to model uncertainty and label consistency compared to standard prompting. Evaluations by human participants further demonstrate that the explanations generated by CLUE are more helpful, more informative, less redundant, and more logically consistent with the input. This work establishes a foundation for explainable fact-checking systems, providing end users with grounded, faithful explanations that reflect the model’s uncertainty. Furthermore, our approach may prove useful for future work in a wide range of information-seeking and retrieval-augmented tasks (e.g., question answering) in which explaining uncertainty given conflicting context is critical.
Limitations
Our paper proposes a novel framework for generating NLEs towards the model’s uncertainty by explicitly pointing to the conflicts or agreements between the claim and multiple pieces of evidence. While our framework demonstrates improved explanation quality through rigorous evaluation across multiple language models and datasets, we acknowledge several limitations.
Our experiments are constrained to medium-sized models (Qwen2.5-14B-Instruct, Gemma2-9B-It, OLMo‑2‑1124‑13B‑Instruct, and DeepSeek‑R1‑Distill‑Qwen‑14B), which are selected based on computational limitations. Although these models show significant improvements over baseline performance, our results suggest that larger models (e.g., 70B parameter scale) may further enhance explanation quality, particularly in terms of coverage and redundancy. Our framework’s modular design readily accommodates such scaling.
CLUE includes an auxiliary relation labeler $L_{\mathrm{rel}}$ that assigns each extracted span interaction one of {agree, disagree, unrelated}. In the main experiments we use GPT-4o for this step to reduce relation-label noise, but this step is optional: $L_{\mathrm{rel}}$ is modular, and a fully open-weight variant (e.g., Qwen2.5-14B-Instruct) yields comparable downstream NLE quality with only a small drop (App. C.2). Reducing this remaining gap while remaining fully open-weight may require improving relation-label prompting (more local context, higher-quality shots) or lightly fine-tuning an open-weight model on a compact relation-judgment set (App. C.4).
Like most multi-stage pipelines, CLUE can occasionally propagate upstream errors: manual inspection reveals failure modes such as fragmentary spans and occasional relation mislabels, which can in turn reduce NLE faithfulness (App. C.3). These issues could be mitigated by encouraging more complete semantic spans during extraction (e.g., mild constraints during community detection), improving relation-label robustness (prompting or lightweight fine-tuning), and adding lightweight generation-time checks that down-weight unreliable span interactions before producing the final explanation (App. C.4).
In this study, we focus on the HealthVer and DRUID datasets, which pair claims with discrete pieces of evidence, making them ideal for studying evidence-conflict scenarios. Our experiments primarily focus on setups where one claim and two or three pieces of evidence are presented. Although we did not explore setups involving more complex claim-evidence interactions, when multiple claims and/or pieces of evidence are introduced, our framework, grounded in pairwise interactions, can seamlessly accommodate more complex scenarios by simply increasing the number of input claim/evidence pieces; no methodological changes are required.
CLUE targets settings where users need to understand why the model is uncertain by grounding uncertainty in concrete claim-evidence and inter-evidence conflicts/agreements. This focus is motivated by qualitative findings that professional fact-checkers want systems to “show the work” behind a decision, especially when evidence-based tensions drive uncertainty, so that users can decide what to verify next and when to trust the output (Warren et al., 2025). While in this paper, we conduct a human study and find that uncertainty explanations generated by our proposed CLUE framework are more helpful than those from the simple prompting method, we do not study whether these uncertainty-oriented explanations are preferable in every context compared to alternative methods for communicating uncertainty (e.g., numeric confidence (Zimmer, 1983; Wallsten et al., 1993; van der Waa et al., 2020; Liu et al., 2020) or verbal hedges (Lin et al., 2022; Mielke et al., 2022; Yona et al., 2024; Kim et al., 2024)). In-situ evaluation with professional fact-checkers, potentially including interface-level ablation studies (e.g., tuning the number of surfaced interactions $k$ and using compact/expandable formats), would be needed to study the effects of uncertainty explanations on efficiency, error detection, and calibrated trust in other contexts.
Our work is limited to the scope of explaining model uncertainty arising from evidence conflicts. While this captures a critical subset of cases, real-world uncertainty may also stem from other sources, including insufficient evidence, knowledge gaps in the model, and context-memory conflicts. We view this work as a foundational step toward broader research on model uncertainty explanation.
Ethical Considerations
This work concerns automated fact-checking, which aims to reduce the harm and spread of misinformation, but nevertheless has the potential for harm or misuse through model inaccuracy, hallucination, or deployment for censorship. Our current work aims to provide explanations that allow users to examine the outputs of these systems more critically, and so we do not see any immediate risks associated with it.
Our work is limited to examining claims, evidence, and explanations in English, and so our results may not be generalisable to other languages. As the task involved complex reasoning about technical subjects, we screened our participants to be native English speakers to ensure that they could fully understand the material and increase the chances of high-quality responses (see I.1 for details). However, this criteria may also introduce or reinforce existing biases and limit the generalisability of our findings. Participants were informed about the study and its aims before agreeing to provide informed consent. No personal data was collected from participants and they received fair payment for their work (approximately 9 GBP/hour).
Acknowledgments
[…]
Appendix A Backbone model performance on public benchmarks
Table 3 summarises the publicly reported five-shot results on two standard reasoning benchmarks. All figures are taken verbatim from the official model cards or accompanying technical reports. Figures are copied from the official model cards.
Model Params MMLU GSM8K
Qwen2.5-14B-Instruct (Qwen Team, 2024) 14.7 B 79.7 90.2
Gemma-2-9B-IT (Gemma Team, 2024) 9.0 B 71.3 68.6
OLMo-2-1124-13B-Instruct (Team OLMo et al., 2025) 13 B 67.5 54.2
Table 3: Benchmark scores on MMLU (Hendrycks et al., 2021) and GSM8K (Cobbe et al., 2021) are used to characterize instruction-following and reasoning strength.
These numbers corroborate our claim that Qwen2.5-14B-Instruct is the strongest of the three for instruction-following and reasoning.
Appendix B Method: Selecting attention heads to steer
Following Zhang et al. (2024b), we steer only a selected subset of attention heads rather than all of them, because targeted steering yields larger gains in output quality. Our selection criterion, however, differs from theirs: instead of ranking heads by their impact on task accuracy, we rank them by how strongly they affect the model’s predictive uncertainty during fact-checking.
Concretely, for each fact-checking dataset chosen in this work(see details in §3.1), $D$, we draw a validation subset $D_{d}$ with $|D_{d}|=300$ examples. For every input $X\in D_{d}$, we compute the model’s baseline uncertainty score $u(X)$ when it predicts the fact-checking label as stated in §2.2. Then, for each attention head identified by layer $\ell$ and index $h$, we zero out that head, re-run the model, and measure the absolute change in uncertainty
$\Delta u(X,\ell,h)\;=\;\bigl|\,u(X)\;-\;u_{/o(l,h)}(X)\bigr|.$
Averaging $\Delta u(X,l,h)$ over all $X\in D_{d}$ yields a single importance score for head $(\ell,h)$. We rank the heads by this score and keep the top $t$ heads for each dataset and each model. Note that we set $t=100$ in line with the recommendation of Zhang et al. (2024b) and to balance steering effectiveness against the risk of degeneration.
Appendix C Conflict and Agreement Identification Details
C.1 Prompt Example for Assigning Relation Labels to Captured Span Interactions
To identify agreements and conflicts between the claim and the two evidence passages, we use the prompt in Figure 3 to label each extracted span interaction (see §2.3).
You are a helpful assistant. Your task:
1. Read the claim and its two evidence passages (E1, E2).
2. For each supplied span interaction, decide whether the two spans
AGREE, DISAGREE, or are UNRELATED, taking the full context into account.
3. Output the span pairs exactly as given, followed by
"relation: agree|disagree|unrelated".
Return format:
1. "SPAN A" - "SPAN B" relation: <agree|disagree|unrelated>
2. ...
3. ...
### SHOT 1 (annotated example)
Claim: [...]
Evidence 1: [...]
Evidence 2: [...]
Span interactions (to be labelled):
1. "[...]" - "[...]"
2. "[...]" - "[...]"
3. "[...]" - "[...]"
Expected output:
1. "[...]" - "[...]" relation: ...
2. "[...]" - "[...]" relation: ...
3. "[...]" - "[...]" relation: ...
### SHOT 2 % omitted for brevity
### SHOT 3 % omitted for brevity
### NEW INSTANCE (pre-filled for each new example)
Claim: {CLAIM}
Evidence 1: {E1}
Evidence 2: {E2}
Span interactions:
1. "{SPAN1-A}" - "{SPAN1-B}"
2. "{SPAN2-A}" - "{SPAN2-B}"
3. "{SPAN3-A}" - "{SPAN3-B}"
Figure 3: Prompt template for span interaction relation labelling.
C.2 Open-weight Alternative for Span-Interaction Relation Labeling
This section studies how the choice of relation labeler affects downstream NLE quality in CLUE. While we use GPT-4o as the default relation labeler to reduce labeling noise, we show that a strong open-weight alternative is also feasible, with only a small performance drop.
As discussed in Sec. 2.3, CLUE uses a relation labeler $L_{\mathrm{rel}}$ to assign each extracted span interaction one of $\{\textsc{agree},\textsc{disagree},\textsc{unrelated}\}$. We considered two labeler options: (i) a closed-API model, which can reduce label noise at the cost of API usage, and (ii) an open-weight model, which is fully local but may yield slightly noisier labels and thus introduce additional noise in downstream NLE generation.
To quantify the impact of the labeler choice, we reran the pipeline with one open-weight alternative for $L_{\mathrm{rel}}$: Qwen2.5-14B-Instruct, keeping all other components fixed. Table 4 reports the results.
HealthVer DRUID
Relation labeler Faith. ($\uparrow$) Span-Cov. ($\uparrow$) Span-Ext. ($\downarrow$) LEE ($\uparrow$) Faith. ($\uparrow$) Span-Cov. ($\uparrow$) Span-Ext. ($\downarrow$) LEE ($\uparrow$)
GPT-4o 0.033 0.44 0.53 0.80 0.102 0.28 0.20 0.77
Qwen2.5-14B-Instruct 0.025 0.41 0.58 0.78 0.076 0.25 0.25 0.75
Table 4: Effect of span-interaction relation labeler choice. We rerun CLUE while swapping only $L_{\mathrm{rel}}$ (GPT-4o vs. Qwen2.5-14B-Instruct) and keeping all other components fixed (backbone: Qwen2.5-14B-Instruct; variant: CLUE-Span+Steering; k=3k{=}3). An open-weight labeler is feasible with only a small drop in downstream NLE quality. Bold indicates the better result per metric within each dataset.
Overall, Qwen2.5-14B-Instruct yields similar, though slightly weaker, downstream NLE quality compared to GPT-4o across both datasets. For instance, on HealthVer, Faithfulness drops from $0.033$ to $0.025$ and LEE from $0.80$ to $0.78$; on DRUID, Faithfulness drops from $0.102$ to $0.076$ and LEE from $0.77$ to $0.75$.
These results indicate that an open-weight labeler is feasible in CLUE, and the framework is not tied to a proprietary component. In this work, we use GPT-4o in the main experiments because it provides consistently stronger relation labeling and reduces the risk of error propagation from noisy relation labels to the final explanations. We further discuss how upstream labeling errors can propagate and mitigation strategies in App. C.3.
A natural concern is whether using a closed-API labeler compromises transparency. We emphasize that CLUE is white-box with respect to the backbone model: uncertainty scoring, span-interaction extraction, and attention steering require access to the backbone model internals, which closed APIs do not provide. Accordingly, running the entire pipeline with GPT-4o would forgo these white-box signals and reduce the method to surface-level prompting; prior work shows such verbalized uncertainty can be unreliable, and fact-checking practice requires systems to “show the work” behind a decision (Yona et al., 2024; Tanneru et al., 2024; Warren et al., 2025).
GPT-4o is used only for the narrow task of pairwise relation judgment after span interactions have already been extracted from backbone internals; as such, it does not affect the transparency of the core framework.
C.3 Error Propagation Analysis
We manually inspected 50 HealthVer instances (top-$3$ span pairs per instance; N=150N{=}150 span pairs) to assess (i) span extraction quality and (ii) relation labeling accuracy when using GPT-4o as $L_{\mathrm{rel}}$.
Span extraction quality.
Most extracted spans are very short: 105/150 span pairs contain two-token spans (often salient nouns such as subjects/objects). In 35/150 cases, at least one span is incomplete (e.g., “D” instead of “vitamin D”) or semantically underspecified (e.g., “is not”). Overall, the extractor tends to favor short, high-salience tokens, suggesting potential for future work to encourage more complete semantic units.
Relation labelling accuracy.
On these 150 span pairs, GPT-4o achieves 86% accuracy (129/150 correct). Among the 21 errors, 12 are cases where an agree relation is mislabelled as unrelated, typically when one span is incomplete (e.g., “D” vs. “vitamin D”), indicating that span completeness is a primary failure mode.
Propagation to downstream NLEs.
Using Qwen2.5-14B-Instruct with CLUE-Span+Steering, 13/21 mislabeled interactions are explicitly mentioned in the generated NLEs, showing that upstream errors can propagate into explanations. Moreover, the subset of instances with mislabeled interactions exhibits lower faithfulness (mean Entropy-CCT 0.020) than the overall HealthVer–Qwen CLUE-Span+Steering average (0.033; Table 1), suggesting that incorrect relation labels can reduce explanation faithfulness.
This analysis suggests that fragmentary spans and relation-label noise can be an upstream error source and carry forward to the final NLEs, motivating mitigation strategies discussed in App. C.4 for future work.
C.4 Mitigating Error Propagation
We outline several directions to mitigate error propagation in future work: (1) Improve span extraction: add mild constraints during community detection so extracted spans more consistently capture complete semantic units rather than clipped tokens. (2) Improve relation labeling robustness: use stronger prompting (broader local context, higher-quality example shots) or lightly fine-tune an open-weight model on a compact relation-judgment set. (3) Generation-time sanity checks: add a lightweight consistency check that flags or down-weights span pairs judged internally inconsistent with the claim/evidence before producing the final NLE.
Appendix D Resources and Model Cards
Appendix E Computational Cost and Resource Requirements
Setup.
We report the computational overhead of CLUE compared to PromptBaseline. Unless otherwise stated, measurements are collected using Qwen2.5-14B-Instruct on the HealthVer dataset with a single NVIDIA A100 80GB GPU. All values are reported per instance (averaged over the evaluation set).
E.1 Latency
Table 5 reports the average wall-clock time per instance for the three methods. Both CLUE-Span and CLUE-Span+Steering are roughly 2 seconds slower per instance than PromptBaseline, and this overhead is entirely due to the additional span-interaction extraction and relation labeling stage.
Approach Uncertainty Scoring (s) Span Extraction + Labeling (s) NLE Generation (s) Total (s)
PromptBaseline 0.364 14.8969 15.2609
CLUE-Span 0.364 2.0452 14.7730 17.1822
CLUE-Span+Steering 0.364 2.0452 14.7874 17.1966
Table 5: Average time cost per instance for the three NLE generation methods on HealthVer using Qwen2.5-14B-Instruct on a single NVIDIA A100 80GB GPU. “–” indicates that span extraction/labeling is not used by PromptBaseline.
E.2 GPU memory usage
Table 6 reports peak GPU memory usage (VRAM). During NLE generation, VRAM usage is similar across all methods. CLUE-Span+Steering requires about 0.32 GiB more VRAM than PromptBaseline, primarily because the prompt is longer (it includes the extracted span interactions; see App. G.2). Notably, CLUE-Span+Steering does not require more VRAM than CLUE-Span, since during generation KV-cache and activations dominate memory and steering does not add a separate memory footprint beyond the base forward pass.
Approach Uncertainty Scoring (GiB) Span Interaction Extraction (GiB) NLE Generation (GiB)
PromptBaseline 29.71 30.52
CLUE-Span 29.71 29.72 30.84
CLUE-Span+Steering 29.71 29.72 30.84
Table 6: Average peak VRAM per instance for the three NLE generation methods (HealthVer, Qwen2.5-14B-Instruct, NVIDIA A100 80GB).
E.3 Memory requirement for steering 100 heads
For selecting the most important 100 attention heads (App. B), the peak VRAM is 29.8 GiB. For steering these heads during NLE generation (CLUE-Span+Steering), the average peak VRAM per instance is 30.84 GiB. Thus, a GPU with 40GB of memory satisfies the requirement for attention steering across 100 heads.
E.4 GPT-4o API cost for relation labeling
When using GPT-4o as $L_{\text{rel}}$, the total relation-labeling cost scales linearly with the number of labeled instances. In our main experiments, the cumulative cost across datasets and backbones was approximately \$30 (about \$0.003 per instance), with an average input length of $\sim$1k tokens and an average output length of $\sim$50 tokens per call.
Appendix F Methodology Supplement: Faithfulness Evaluation Details
This supplement provides implementation and statistical details for the Entropy-CCT faithfulness metric introduced in §4.1. We describe (i) how counterfactual perturbations are generated, (ii) how Entropy-CCT relates to the original CCT, and (iii) how we test statistical significance and report results.
F.1 Entropy-CCT Perturbation Generation
To evaluate how faithfully each NLE reflects model uncertainty, we generate multiple counterfactuals per instance, following Atanasova et al. (2020) and Siegel et al. (2024) (see §4.1). For every input, comprising one claim and two evidence passages, we first tag part‐of‐speech with spaCy (Honnibal and Montani, 2017) (en_core_web_sm), then choose four random insertion sites (uniformly among tokens tagged as NOUN or VERB). At each site we insert either (i) a random adjective before a noun or (ii) a random adverb before a verb. The candidate modifiers are drawn uniformly from WordNet (Miller, 1992) using NLTK (Bird et al., 2009): we construct adjective/adverb candidate lists by collecting all lemma names from all WordNet adjective/adverb synsets, respectively. Because we sample three random candidates for each of the four positions, this procedure yields $4\times 3=12$ perturbations per instance, providing a sufficient set for the subsequent Entropy-CCT evaluation, in which we check whether the NLE mentions the inserted word and compute the correlation between that mention and the uncertainty change induced by each perturbation.
F.2 Entropy‑CCT vs. Original CCT
CCT Siegel et al. (2024) evaluates faithfulness by perturbing the input and measuring the resulting change in the model’s predicted label distribution. Specifically, it uses the Total Variation Distance (TVD) between the pre- and post-perturbation distributions $P(\cdot\mid X)$ and $P(\cdot\mid X^{\prime})$:
$\mathrm{TVD}(P,Q)=\frac{1}{2}\sum_{i}|P_{i}-Q_{i}|.$ (12)
Entropy-CCT is a minimal adaptation of CCT that preserves the same perturbation procedure and the same correlational test: we correlate the perturbation impact with whether the inserted token is mentioned in the NLE. The only change is the signal we correlate. Instead of TVD (probability-shift), we use Absolute Entropy Change (AEC), i.e., the change in predictive entropy, to directly target uncertainty (our object of explanation).
F.3 Statistical Testing & Significance
We test statistical significance of $r_{\text{pb}}$ using the standard Pearson correlation $t$-test. Each perturbation constitutes one observation: for each dataset we evaluate 600 instances with 12 perturbations per instance, yielding $n=600\times 12=7{,}200$ observations (App. F.1). Because the point-biserial correlation is algebraically equivalent to Pearson correlation between a continuous and a binary variable, we use the standard Pearson $t$-test for $H_{0}:r_{\text{pb}}=0$.
Interpreting $r_{\text{pb}}$ and $\Delta r_{\text{pb}}$. The Entropy-CCT score is the point-biserial correlation (Tate, 1954) between the absolute entropy change $|\Delta u|$ and the binary mention flag $m$. Because it is mathematically identical to a Pearson $r$ computed between one continuous and one binary variable, it obeys $1\leq r_{\text{pb}}\leq 1$. When $r_{\text{pb}}=0$, it means the high- and low-impact perturbations are mentioned equally often. If the two strata are roughly balanced, every +0.01+0.01 in $r_{\text{pb}}$ increases the probability that a truly uncertainty-influential token is mentioned by about one percentage point (pp). A gain $\Delta r_{\text{pb}}$ therefore translates to an absolute improvement of $\approx|\Delta r_{\text{pb}}|\times 100$ pp in mention rate. For instance, moving from 0.08-0.08 to +0.06+0.06 is a swing of $0.14$, corresponding to about $14$ pp.
Significance testing. Because the point-biserial is a Pearson correlation, the familiar tt–test applies:
$t$ $=r_{\text{pb}}\,\sqrt{\frac{n-2}{1-r_{\text{pb}}^{2}}},$ (13)
$t$ $\sim t_{(n-2)}\qquad\text{under }H_{0}\!:r_{\text{pb}}=0.$ (14)
With $n=7,200$ we have $\text{df}=7,198$; the critical two-sided values are $|t|>1.96$ for $p<0.05$ and $|t|>2.58$ for $p<0.01$.
Faithfulness with significance results. Table  7 shows the point-biserial coefficients $r_{\text{pb}}$, which is our faithfulness measurement for model uncertainty(See, Eq.9), the associated $t$ statistics, and two-sided $p$ values for every model–method pair. Values that meet the stricter $p<0.01$ criterion are highlighted in bold.
Model Method $r_{\text{pb}}$ $t$ $p$
HealthVer
Qwen2.5-14B-Instruct PromptBaseline 0.028-0.028 2.38-2.38 $1.7\times 10^{-2}$
CLUE-Span +0.006+0.006 +0.51+0.51 $6.1\times 10^{-1}$
CLUE-Span+Steering +0.033+0.033 +2.80+2.80 $\mathbf{5.1\times 10^{-3}}$
OLMo-2-1124-13B-Instruct PromptBaseline 0.100-0.100 8.53-8.53 $\mathbf{<10^{-15}}$
CLUE-Span +0.005+0.005 +0.42+0.42 $6.7\times 10^{-1}$
CLUE-Span+Steering +0.020+0.020 +1.70+1.70 $9.0\times 10^{-2}$
Gemma-2-9B-IT PromptBaseline 0.105-0.105 8.96-8.96 $\mathbf{<10^{-15}}$
CLUE-Span +0.007+0.007 +0.59+0.59 $5.5\times 10^{-1}$
CLUE-Span+Steering +0.021+0.021 +1.78+1.78 $7.5\times 10^{-2}$
DRUID
Qwen2.5-14B-Instruct PromptBaseline 0.080-0.080 6.81-6.81 $\mathbf{9.8\times 10^{-12}}$
CLUE-Span +0.089+0.089 +7.58+7.58 $\mathbf{3.4\times 10^{-14}}$
CLUE-Span+Steering +0.102+0.102 +8.70+8.70 $\mathbf{<10^{-15}}$
OLMo-2-1124-13B-Instruct PromptBaseline 0.130-0.130 11.12-11.12 $\mathbf{<10^{-15}}$
CLUE-Span +0.014+0.014 +1.19+1.19 $2.3\times 10^{-1}$
CLUE-Span+Steering +0.099+0.099 +8.44+8.44 $\mathbf{<10^{-15}}$
Gemma-2-9B-IT PromptBaseline 0.120-0.120 10.26-10.26 $\mathbf{<10^{-15}}$
CLUE-Span +0.043+0.043 +3.65+3.65 $\mathbf{2.6\times 10^{-4}}$
CLUE-Span+Steering +0.098+0.098 +8.35+8.35 $\mathbf{<10^{-15}}$
Table 7: Detailed faithfulness evaluation results for baseline method PromptBaseline, and two variants of our CLUE framework CLUE-Span and CLUE-Span+Steering on Healthver and Druid dataset based on Qwen2.5-14B-Instruct(Qwen Team (2024)), OLMo-2-1124-13B-Instruct(Team OLMo et al. (2025))and Gemma-2-9B-IT(Gemma Team (2024)). Point-biserial correlation $r_{\text{pb}}$ is our Entropy-CCT measurement(§4.1), along with $t$ statistic and two-sided $p$-value for each model–method pair (n=7,200n=7{,}200, df=7,198df=7{,}198). Entries with $p<0.01$ are bold.
Across both datasets and all three backbones, the PromptBaseline exhibits negative correlations, implying a non-faithful tendency to highlight low-impact tokens in the generated NLEs, with $\text{mean}=-0.094$. The prompt-only variant of our CLUE framework CLUE-Span neutralises this bias and turns the average into +0.027+0.027; two of its coefficients are clear $p<0.01$, indicating a modest but significant improvement regarding faithfulness.
The full CLUE-Span+Steering variant pushes the mean to +0.062+0.062 and achieves $p<0.01$ in four of six settings. Interpreting these numbers using the discussion above, the switch from 0.094-0.094 to +0.062+0.062 yields an absolute increase of $(0.062-(-0.094))\times 100!\approx 16$, pp in the probability that a truly uncertainty-influential token is named in the NLE, which is easily noticeable in qualitative inspection.
The consistently positive, statistically significant gains therefore substantiate the claim made in the main text: CLUE produces markedly more faithful NLEs towards model uncertainty than the PromptBaseline, and the steer variant is particularly beneficial for models that initially struggle with uncertainty attribution.
Appendix G Prompt template for PromptBaseline, CLUE-Span and CLUE-Span+Steering on Healthver and Druid dataset
We designed two prompt templates for our experiments. The baseline prompt (Figure 4) gives the model no span interactions; instead, it must first identify the relevant agreements or conflicts and then discuss them in its explanation. In contrast, the prompt used by our CLUE framework (Figure 5) supplies the three pre-extracted span interactions (§2.3). The model is explicitly instructed to base its explanation on these spans, ensuring that the rationale remains grounded in the provided evidence.
G.1 Prompt template for PromptBaseline
To generate NLEs about model uncertainty without span-interaction guidance, we craft a three-shot prompt that instructs the model to identify the interactions most likely to affect its uncertainty and to explain how these relations they represent affect it. (See Figure 4).
You are a helpful assistant. Your tasks:
1. Determine the relationship between the claim and the two evidence passages.
2. Explain your predictions uncertainty by identifying the three most
influential span interactions from Claim-Evidence 1, Claim-Evidence 2,
and Evidence 1-Evidence 2, and describing how each interactions relation
(agree, disagree, or unrelated) affects your overall confidence.
Return format: [Prediction] [Explanation]
### SHOT 1
Input
Claim: [...]
Evidence 1: [...]
Evidence 2: [...]
Output
[Prediction: ...] [Explanation: ...]
### SHOT 2 % omitted for brevity
### SHOT 3 % omitted for brevity
### NEW INSTANCE
Claim: {CLAIM}
Evidence 1: {E1}
Evidence 2: {E2}
Your answer:
Figure 4: Three-shot prompt for PromptBaseline (Shots 2-3 omitted) on the HealthVer and DRUID datasets.
G.2 Prompt template for CLUE-Span and CLUE-Span+Steering
To generate NLEs about model uncertainty with the span-interaction guidance, we craft a three-shot prompt that instructs the model to discuss how these interactions, along with the relations they represent, affect its uncertainty. (See Figure 5).
You are a helpful assistant. Your tasks:
1. Determine the relationship between the claim and the two evidence passages.
2. Explain your predictions uncertainty by referring to the three span
interactions provided below (Claim-Evidence 1, Claim-Evidence 2,
Evidence 1-Evidence 2) and describing how each interactions relation
(agree, disagree, or unrelated) affects your overall confidence.
Return format: [Prediction] [Explanation]
### SHOT 1
Input:
Claim: [...]
Evidence 1: [...]
Evidence 2: [...]
Span interactions:
1. ’’[...]’’ - ’’[...]’’ (C-E1) relation: [...]
2. ’’[...]’’ - ’’[...]’’ (C-E2) relation: [...]
3. ’’[...]’’ - ’’[...]’’ (E1-E2) relation: [...]
Output:
[Prediction: ...] [Explanation: ...]
### SHOT 2 % omitted for brevity
### SHOT 3 % omitted for brevity
### NEW INSTANCE
Claim: {CLAIM}
Evidence 1: {E1}
Evidence 2: {E2}
Span interactions (pre-filled):
1. ’’{SPAN1-A}’’ - ’’{SPAN1-B}’’ (C-E1) relation: {REL1}
2. ’’{SPAN2-A}’’ - ’’{SPAN2-B}’’ (C-E2) relation: {REL2}
3. ’’{SPAN3-A}’’ - ’’{SPAN3-B}’’ (E1-E2) relation: {REL3}
Your answer:
Figure 5: Three-shot prompt for CLUE-Span and CLUE-Span+Steering (Shots 2-3 omitted) on the HealthVer and DRUID datasets.
Appendix H Extended Evaluation Results
H.1 Extended Evaluation Results for Scenarios Involving Three Evidence pieces
Metrics PromptBase CLUE-S CLUE-Steer
Faith. ($\uparrow$) 0.08 0.10 0.13
Span-Cov. ($\uparrow$) 0.25 0.41
Span-Ext. ($\downarrow$) 0.23 0.27
LEE ($\uparrow$) 0.80 0.86 0.86
Table 8: Automatic evaluation for three types of NLE on HealthVer using Qwen2.5-14B-Instruct when one claim and three pieces of evidence are involved: PromptBase(PromptBaseline) is the baseline method, CLUE-S (CLUE-Span) and CLUE-Steer(CLUE-Span+Steering) are the two variants of our CLUE framework.
To verify the generalizability of our framework in scenarios involving multiple evidence pieces, we conduct an experiment on the Healthver dataset using the Qwen2.5‑14B‑Instruct model, pairing each claim with three pieces of evidence instead of two (as shown in Tab. 8). Note that we select the top‑$k=6$ span interactions for NLE generation, taking the top‑1 interaction from each pairwise combination among the $C$, $E_{1}$, $E_{2}$, and $E_{3}$ input parts.
Overall, we observe a similar trend with three pieces of evidence as with two (see §4.5): CLUE-Span+Steering and CLUE-Span show better Faithfulness and Label-Explanation Entailment than PromptBaseline. Interestingly, the Faithfulness of all three explanation types increases relative to the two‑evidence setting, from -0.028 to 0.08 for PromptBaseline, from 0.006 to 0.10 for CLUE-Span, and from 0.033 to 0.13 for CLUE-Span+Steering. This may be because, when NLEs are grounded in richer evidence, they capture more nuances in both claim-evidence and inter‑evidence interactions, thereby more faithfully reflecting the model’s fact‑checking decision.
It is notable that CLUE-Span+Steering includes more targeted span interactions (higher Span‑Coverage: 0.41 > 0.27) as well as more redundant span interactions (higher Span‑Extraneous: 0.27 > 0.23) than CLUE-Span. This suggests that, although CLUE-Span+Steering better focuses on target interactions, it still suffers from redundancy due to the inclusion of more irrelevant interactions, which we call for future work to address.
H.2 Extended Evaluation Results on Reasoning Model
To demonstrate that our framework generalizes to reasoning models trained primarily via large‑scale reinforcement learning, and designed to encourage chain‑of‑thought exploration, we adopt DeepSeek‑R1‑Distill‑Qwen‑14B [^5] https://huggingface.co/deepseek-ai/DeepSeek-R1-Distill-Qwen-14B (DeepSeek-AI, ) as the base model and evaluate it on the HealthVer dataset in the setting with one claim and two pieces of evidence (see Tab. 9).
Metrics PromptBase CLUE-S CLUE-Steer
Faith. ($\uparrow$) 0.12 0.16 0.18
Span-Cov. ($\uparrow$) 0.58 0.67
Span-Ext. ($\downarrow$) 0.59 0.64
LEE ($\uparrow$) 0.88 0.90 0.89
Table 9: Automatic evaluation for three types of NLE on HealthVer using DeepSeek-R1-Distill-Qwen-14B when one claim and two pieces of evidence are involved: PromptBase(PromptBaseline) is the baseline method, CLUE-S (CLUE-Span) and CLUE-Steer(CLUE-Span+Steering) are the two variants of our CLUE framework.
Overall, both variants of our CLUE model, CLUE-Span and CLUE-Span+Steering, show higher Faithfulness and Label‑Explanation Entailment than the baseline method PromptBaseline, confirming the effectiveness of our framework. It is also notable that, compared with the evaluation results on other models in Tab. 1, all three types of explanations generated with the DeepSeek‑R1‑Distill‑Qwen‑14B model achieve the best performance across metrics, especially the highest Faithfulness and Span‑Coverage. The former implies that reinforcement-learning-trained reasoning models may be more capable of generating highly faithful explanations with the thinking step; the latter suggests that this type of model follows our interaction-guided instructions more effectively. Consequently, both CLUE-Span and CLUE-Span+Steering include more target-span interactions, yielding higher Span-Coverage.
Comparing the two variants, CLUE-Span+Steering exhibits higher Span‑Coverage (0.67 > 0.58) than CLUE-Span, but also more Span‑Extraneous (0.64 > 0.59), suggesting that CLUE-Span+Steering tends to include redundant interactions while focusing on target interactions. CLUE-Span+Steering also achieves the highest Faithfulness (0.18 > 0.16), consistent with its higher Span‑Coverage, but yields similar Label‑Explanation Entailment to CLUE-Span (0.89 vs. 0.90), likely due to the increased Span‑Extraneous. Addressing this trade‑off is left for future work.
Appendix I Human Evaluation Details
I.1 Participants and Materials
Participants
We recruited N=12 participants from Prolific, screened to be native English speakers from Australia, Canada, Ireland, New Zealand, the United Kingdom, and the United States. The study was approved by our institution’s Research Ethics Committee (reference number anonymised).
Materials
Forty claims (20 from DRUID, 20 from HealthVer) were selected at random for evaluation, while 120 explanations were evaluated in total. For each instance, participants were provided with a claim, two evidence documents, model verdict, model numerical certainty, and three alternative explanations (see Figure 6 in I.6). The explanations presented to participants were those generated using Qwen2.5-14b-instruct Qwen Team (2024) based on its automatic evaluation performance.
Each participant evaluated explanations for 10 instances (5 labelled ‘True’, 5 labelled ‘False’), in addition to two attention check instances which were used to screen responses for quality. To mitigate the difficulty and cognitive demands of the task, we kept the number of instances per participant low, and removed participants who failed attention checks from the data. In order to minimise the risk of bias, each explanation presented to participants was assigned a neutral label: ‘Explanation A’, ‘Explanation B’, ‘Explanation C’, so that participants could not infer anything about how the explanations were generated or indicate that one should be favoured over another. Each explanation was also presented in a different, random order for each instance (e.g., ABC for Claim 1, CAB for Claim 2, etc.) to mitigate order effects such as primacy bias or recency bias. The order in which the response options ‘Explanation A’, ‘Explanation B’, ‘Explanation C’ also appeared in the same random order.
Procedure
Participants read information about the study (see I.3) and provided informed consent (see I.4) before reading detailed task instructions and completing a practice example of the task (see I.5). Participants then progressed through the study at their own pace. The task took approximately 20 minutes, and participants were paid £3 for their work.
I.2 Human Evaluation Results
I.2.1 Interrater agreement
In line with similar NLE evaluations carried out by previous studies (e.g., Atanasova et al. (2020)), interrater agreement (Kendall’s W Kendall and Smith (1939)) was moderate to low (see Table 10). We attribute this to the relative complexity of the task and individual differences in how the information was perceived.
DRUID HealthVer
Set A Set B Set A Set B
Helpfulness .016 .079 .003 .013
Consistency .44 .058 .017 .016
Non-redundancy .005 .084 .005 .019
Coverage .494 .113 .018 .027
Overall Quality .005 .158 .01 .002
Table 10: Interrater agreement (Kendall’s W) for human evaluation
I.2.2 CLUE Variant preferences
PromptBase CLUE-S CLUE-Steer
Helpfulness
Overall 2.025 1.892 1.867
DRUID 1.9 1.917 1.767
HealthVer 2.15 1.867 1.967
Consistency
Overall 1.875 1.783 1.817
DRUID 1.717 1.75 1.617
HealthVer 2.033 1.817 2.017
Non-redundancy
Overall 2.05 1.908 1.833
DRUID 1.983 1.983 1.683
HealthVer 2.117 1.833 1.983
Coverage
Overall 1.967 1.775 1.758
DRUID 1.767 1.75 1.617
HealthVer 2.167 1.8 1.9
Overall Quality
Overall 1.967 1.908 1.925
DRUID 1.9 1.9 1.817
HealthVer 2.033 1.917 2.033
Table 11: Mean Average Rank (MAR) for the five human-evaluation criteria applied to explanations from Qwen2.5-14B-Instruct on the HealthVer and DRUID datasets (chosen for its high faithfulness; see §4.5). PromptBaseline, CLUE-Span (CLUE-S), and CLUE-Span+Steering (CLUE-Steer) are compared. Lower MAR means a better (higher) average rank; the best score in each row is boldfaced.
Table 11 shows the mean rank assigned by participants to each explanation variant. The explanations generated by CLUE were preferred by our participants to those generated using PromptBaseline: the explanations generated by CLUE-Span+Steering were rated as most helpful, highest coverage, and containing the least amount of redundant information, while those from CLUE-Span were judged to have the highest consistency and overall quality.
Finally, we observed slight variation between datasets: CLUE-Span+Steering tended to be rated higher than CLUE-Span for DRUID, and vice versa for HealthVer. This may arise from differences in length and complexity of the input: DRUID evidence documents, retrieved from heterogeneous online sources and often consisting of longer form new articles, may have benefited from attention steering more than HealthVer evidence documents which consist of focused, shorter extracts from scientific abstracts.
I.3 Human Evaluation Information Screen
Thank you for volunteering to participate in this study! Before you decide whether you wish to take part, please read this information screen carefully.
1. What is the project about?
Our goal is to make sure that AI fact-checking systems can explain the decisions they produce in ways that are understandable and useful to people. This survey is part of a project to help us understand [anonymised]
2. What does participation entail?
You are invited to help us explore what kinds of explanations work better in fact-checking. In this task you will see claims, an AI system’s prediction about whether this claim is true or false and corresponding evidence used to make the prediction. You will also see an explanation for why the AI system is certain or uncertain about its prediction to help you decide how to interpret the true/false prediction. We ask you to evaluate the explanations along 5 different dimensions (the detailed explanation of the task is on the next page). All participants who complete the survey will receive a payment of £3. There is no cost to you for participating. You may refuse to participate or discontinue your involvement at any time without penalty.
3. Source of funding
This project has received funding from [anonymised]
4. Consenting to participate in the project and withdrawing from the research
You can consent to participating in this study by ticking the box on the next page of the study. Participation in the study is completely voluntary. Your decision not to consent will have no adverse consequences. Should you wish to withdraw during the experiment you can simply quit the webpage. All incomplete responses will be deleted. After you have completed the study and submitted your responses, it will no longer be possible to withdraw from the study, as your data will not be identifiable and able to linked to you.
5. Possible benefits and risks to participants
By participating in this study you will be contributing to research related to understanding what kinds of explanations are useful to people who use or who are impacted by automated fact checking systems. This is a long-term research project, so the benefits of the research may not be seen for several years. It is not expected that taking part will cause any risk, inconvenience or discomfort to you or others.
6. What personal data does the project process?
The project does not process any personal data.
7. Participants’ rights under the General Data Protection Regulation (GDPR)
As a participant in a research project, you have a number of rights under the GDPR. Your rights are specified in [anonymised]
8. Person responsible for storing and processing of data
[anonymised]
Please click ’Next’ to read more about consenting to participate in the study.
I.4 Human Evaluation Consent Form
We hereby request your consent for processing your data. We do so in compliance with the General Data Protection Regulation (GDPR). See the information sheet on the previous screen for more details about the project and the processing of your data.
  • I confirm that I have read the information sheet and that this forms the basis on which I consent to the processing of my data by the project.
  • I hereby give my consent that [anonymised] may register and process my data
  • I understand that any data I provide will be anonymous and not identifiable to me.
  • I understand that my anonymous response data will be retained by the study team.
  • I understand that after I submit my responses at the end of the study, they cannot be destroyed, withdrawn, or recalled, because they cannot be linked with me.
  • I understand that there are no direct benefits to me from participating in this study
  • I understand that anonymous data shared through publications or presentations will be accessible to researchers and members of the public anywhere in the world, not just the EU.
  • I give my consent that the anonymous data I provided may be stored in a database for new research projects after the end of this project.
  • I give permission for my anonymous data to be stored for possible future research related to the current study without further consent being required.
  • I understand I will not be paid for any future use of my data or products derived from it.
By checking this box, I confirm that I agree to the above and consent to take part in this study.
$\Box$ I consent
I.5 Evaluation Task Instructions
What do I have to do?
In this study you will see claims, an AI system’s prediction about whether this claim is true or false, how certain the system is about its label, and the corresponding evidence used to make the prediction. You will also see three different explanations for why the AI system is certain or uncertain about its prediction. These explanations are intended to help you decide how to interpret the true/false prediction.
Your task is to evaluate the quality of the explanations provided, not the credibility of the claims and evidence.
What information will I be shown?
You will be shown examples of claims, evidence document, verdicts and explanations.
  • A claim is some statement about the world. It may be true, false, or somewhere in between.
  • Additional information is typically necessary to verify the truthfulness of a claim - this is referred to as evidence or evidence document. An evidence document consists of one or several sentences extracted from an external source for the particular claim. In this study, you will see two evidence documents that have been retrieved for a claim. These evidence documents may or may not agree with each other.
  • Based on the available evidence, a verdict is reached regarding whether a claim is true or false.
  • Uncertainty often arises when evaluating the claim and evidence to reach a verdict. Each verdict is accompanied by a numerical uncertainty score which represents the AI system’s confidence that its predicted verdict is correct.
  • You will see 3 alternative explanations for where uncertainty arises with regard to the verdict. Note that these explanations focus on the AI system’s uncertainty, not the verdict itself.
  • You are asked to evaluate the explanations according to 5 different properties. The properties are as follows:
    Helpfulness. The explanation contains information that is helpful for evaluating the claim and the fact check.
    Coverage. The explanation contains important, salient information and does not miss any important points that contribute to the fact check.
    Non-redundancy. The explanation does not contain any information that is redundant/repeated/not relevant to the claim and the fact check.
    Consistency. The explanation does not contain any pieces of information that are contradictory to the claim and the fact check.
    Overall Quality. Rank the explanations by their overall quality.
  • Please rank the explanations in descending order. For example, you should rank the explanation that you think is most helpful as ‘1’, and the explanation that you think is least helpful as ‘3’. If two explanations appear almost identical, you can assign them the same ranking, but as a general rule, you should try rank them in hierarchical order.
  • The three explanations, Explanation A, Explanation B, and Explanation C, will appear in a different order throughout the study, so you may need to pay some attention to which is which.
Important: Please only consider the provided information (claim, evidence documents, and explanations) when evaluating explanations. Sometimes you will be familiar with the claim, but we ask you to approach each claim as new, whether or not you have seen it before. It doesn’t matter whether you personally agree or disagree with the claim or evidence - we are asking you to evaluate what the AI produces: if you were to see this claim for the first time, would you find the explanation provided by the AI useful? On the next page, you will see an example of the task.
I.6 Example of human evaluation set-up
Here is an example of what you will see during the study. First, you will see a Claim, and two pieces of Evidence, along with an AI system’s predicted Verdict and the system’s Certainty that its prediction is correct.
The parts of the claim and evidence that are most important to the AI system’s certainty are highlighted. Parts of the Claim are Red, parts of Evidence 1 are Blue, and parts of Evidence 2 are Green.
Underneath, you will see three alternative explanations for the AI system’s certainty, Explanation A, Explanation B, and Explanation C. The parts of each explanation that refer to the claim and evidence are colour coded in the same way (Claim = Red, Evidence 1 = Blue, Evidence 3 = Green).
Your task is to read the claim, evidence, and explanations, and rank each explanation based on five properties.
Now, you can try this example below!
Figure 6: Example of human evaluation set-up. Explanation A was generated using PromptBaseline, Explanation B by CLUE-Span, and Explanation C by CLUE-Span+Steering